1.6 g of Ca and 2.60 g of Zn when treated with an acid in excess separately, produced the same amount of hydrogen. If the equivalent weight of Zn is 32.6, what is the equivalent weight of Ca?
A) 10.          B) 20.          C) 40.         D) 5 .  

Grade:9

2 Answers

Arun
5 years ago
Let’s assume Zn and Ca is treated with Hydrochloric acid.
Zn + 2HCl = ZnCl2 + H2
According to reaction, 32.6 g of Zn gives 2 g of Hydrogen gas.
That means 2.6 g of Zn would give 2 X 2.6/32.6 = 0.16 g of Hydrogen gas
Same amount of hydrogen gas (0.16 g) is produced from Ca when treated with acid (HCl)
Ca + 2HCl = CaCl2 + H2
According to reaction, 1 equivalent of Ca would give 2 g of hydrogen gas.
And we know that, 1.6 g of Ca gives 0.16 g of hydrogen gas.
Then, the equivalent weight of Ca is 1.6 X 0.16\2 = 20
Equivalent weight of Ca is 20.
 
Prakhar Upadhyay
5 years ago
Will you please explain how did you come to the conclusion that 32.6 g of Zn gives 2 g of hydrogen gas?